# CGCOD: Class-Guided Camouflaged Object Detection

### v1

@@ 4.2 Class-Guided Detector (CGD)
Overview. To improve the performance of COD, we further propose a novel detector CGD. Which leverages class-promt information to accurately segment camouflaged objects and suppress background noise. Given an input $I_{r}$ of size $W\times H$, We first use PVTv2 Wang et al. (2021) as the visual backbone to extract a set of depth features $\{X_{i}\}_{i=0}^{4}$. As shown in Fig. 3, the semantic-rich feature $f_{4}$, the feature $F_{v}$, and the initially fused multi-modal information $F_{m}$ are input into the class-semantic guidance (CSG). The module aims to explore the internal correlations between the backbone features and the other modalities from both shallow and deep layers, thus generating class-semantic guidance map $G_{c}$ that embed both class semantics and camouflaged semantics. These features $\{X_{i}\}_{i=1}^{4}$ and $G_{c}$ are then passed to the semantics consistency module to further extract fine-grained features related to the camouflaged object, ultimately leading to an accurate segmentation. The process iteratively accumulates intermediate segmentation results to decouple the final output $\{P_{i}\}_{i=1}^{4}$.
@@
Semantics Consistency Module (SCM). Traditional COD models rely on a visual backbone that extracts camouflaged semantics containing both foreground and background clues, often resulting in false negatives and false positives. To address these issues, we propose the SCM. By class guided features, the SCM pays more attention to features with high correlation with guided features, alleviating the interference of irrelevant noise information. Most camouflage is achieved by using channel information (such as color) and spatial location in relatively few cases. So the SCM consists of two stages: the first stage, spatial localization, focuses on capturing spatial information of visual features related to guidance feature, while the second stage, channel refinement, explores inter-channel relationships and adaptively selects features that align with class semantics to suppress background interference. Specifically, we utilize two parallel branches to process multi-level features $X_{i}$ and $X_{i+1}$ independently, facilitating their interaction with class semantic knowledge. Each branch incorporates weighted summation and dot product operations, ensuring a comprehensive interaction between the object class semantic features and the multi-level features. This interaction establishes an initial alignment between the camouflaged foreground and background semantics and the class-guided semantic features. Consequently, it produces a new class-semantic relevance matrix $G^{\prime}_{\text{c}}$ and enriched class semantic features $\tilde{X}_{i+1}$ and $\tilde{X}_{i}$. The process is formulated as follows:
$R_{\text{cls}}=\text{Softmax}\left(\text{Conv}\big{(}X_{i}\oplus X_{i+1}\big{)}\otimes G_{\text{c}}\right)$ (6)
$\hat{X}_{i}=R_{\text{cls}}\otimes X_{i},\quad\hat{X}_{i+1}=R_{\text{cls}}\otimes X_{i+1}$ (7)
$R^{\prime}_{\text{cls}}=\text{FC}\left(\text{Concat}\big{(}\hat{X}_{i+1},\hat{X}_{i}\big{)}\right)$ (8)
$\tilde{X}_{i}=R^{\prime}_{\text{cls}}\oplus\hat{X}_{i},\quad\tilde{X}_{i+1}=R^{\prime}_{\text{cls}}\oplus\hat{X}_{i+1}$ (9)
@@
Here, $i$ denotes the $i$-th operator, $\text{FC}()$ refers to linear projection, while $\text{Conv}()$ signifies a $3\times 3$ convolution.
@@
FSEL Sun et al. (2024) None PVT 0.885 0.857 0.040 0.942 0.877 0.799 0.021 0.937 0.892 0.852 0.030 0.941 - - - -
RISNet Wang et al. (2024) Depth Map PVT 0.870 0.827 0.050 0.922 0.873 0.799 0.025 0.931 0.882 0.834 0.037 0.926 - - - -
VSCodes Luo et al. (2024) 2D Prompt Swin 0.873 0.820 0.046 0.925 0.869 0.780 0.025 0.931 0.882 0.841 0.032 0.935 - - - -
@@
CGNet (Ours) CamoClass CLIP, PVT 0.894 0.863 0.036 0.946 0.890 0.824 0.018 0.948 0.903 0.868 0.027 0.948 0.931 0.902 0.017 0.972
To enhance the feature expression and focus on class-relevant information, we propose a channel-adaptive refinement mechanism. Initially, features $\tilde{X}_{i+1}$ and $\tilde{X}_{i}$ are concatenated along the channel dimension to form a unified feature representation $X_{\text{cot}}$. This representation is then partitioned into four subgroups $\{g_{1},g_{2},g_{3},g_{4}\}$ along the channel dimension. For each subgroup $g_{i}$, its concern weight $w_{i}$ is calculated through the linear layer and sigmoid activation function. These attention weights modulate the importance of each subgroup by element-wise multiplication, emphasizing features that are consistent with the semantics of the object class. The weighted subgroups are subsequently fused along the channel dimension to produce the refined feature representation $F_{\text{s}}$, which highlights the critical information associated with the target class. The process is defined as follows:
$X_{\text{cot}}=\text{Conv}(\text{Concat}(\tilde{X}_{i},\tilde{X}_{i+1}))$ (10)
$\{g_{1},g_{2},\dots,g_{n}\}=\text{Split}(X_{\text{cot}})$ (11)
$w_{i}=\text{Gate}(g_{i})$ (12)
$g^{\prime}_{i}=\text{Conv}_{i}(x_{\text{cat}}\otimes w_{i})$ (13)
$F^{i}_{\text{s}}=\text{Concat}(g^{\prime}_{1},g^{\prime}_{2},\dots,g^{\prime}_{n})$ (14)
Here, Gate()()( ) refers to tow linear projection and a sigmoid, while Conv()()( ) signifies a $3\times 3$ convolution.
@@ 4.3 Loss Function
Since our ultimate goal is to generate a binary foreground graph, a structured loss function consisting of BCE loss and IoU loss can be used in the optimization process Wei et al. (2020). Specifically, the loss function can be expressed as:
$\mathcal{L}(P,G)=\sum_{i=1}^{6}\mathcal{L}_{\text{bce}}(P_{i},G)+\mathcal{L}_{\text{iou}}(P_{i},G),$ (15)
where $P=\{F_{m},F_{v},P_{4},P_{3},P_{2},P_{1}\}$ represent the predictions generated by our model, and $G$ refers to the ground truth annotation.
@@ 5.1 Implementation Details
Data Settings. We used the CamoClass dataset to evaluate our approach. The training set consists of CAMO Le et al. (2019) and COD10K Fan et al. (2022), with a total of 4040 samples and their class-level textual information. The test set consists of test samples from four benchmark datasets, totaling 8615 samples and their class-level textual information. It is important to note that we have fine-grained segmentation of the multi-class samples in the benchmark dataset, so that our total number of samples exceeds the original number of these four benchmark datasets.
Evaluation Metrics. In this work, we use five widely recognized and standard metrics to assess the performance of our method, including S-measure ($S_{\alpha}$an et al. (2017), E-measure ($E^{m}_{\phi}$an et al. (2018), weighted F-measure ($F_{\beta}^{\omega}$v et al. (2021), average F-measure ($F^{m}_{\beta}$v et al. (2021) and mean absolute error ($M$Perazzi et al. (2012).
Training Settings. Following the setup in previous studies Xu et al. (2023, 2022), the parameters of the CLIP model were frozen during the training process, while the parameters of the remaining network components were learnable. We use the Adam optimizer with as learning rate of $1\times 10^{-4}$ to optimize the network parameters, with a batch size of 10 and a total of 200 epochs. We apply various data augmentation techniques to the training data, including random cropping, horizontal flipping, and color jittering. Specially, during both the training and inference phases, the image input size for the CLIP model was $336\times 336$ pixels, while the input size for the class-guided detector was $448\times 448$ pixels. All experiments are conducted on a server equipped with an NVIDIA RTX 3090 (with 24GB memory).
@@ 5.2 Comparison with State-of-the-Art Methods
Quantitative Evaluation. As shown in Tab. 1, CGNet consistently surpasses competing methods across all datasets, showing a significant and substantial performance advantage. And CGNet are on average 3.5% and 5.5% better than the second best method. Due to CRIS, CGFormer, and ASDA are not specifically designed for the COD task, their performance on the four datasets is poor. Moreover, we also present the results of detectors optimized under the class prompt generator (CPG) and proposed CamoClass. In Tab. 1, CPG improve other detectors by 19.73% (SINet), 13.3% (BGNet), 12.68% (PFNet), 4.43% (FPSNet), and increase our class-guided detector by 6.24% (PVT), verifying that CPG serve as effective plug-and-play module for performance improvement.
Table 2: Comparison of methods on seen and unseen classes ($\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$), the best results are marked in red.
Model Seen (2212 images) Unseen (4261 images)
$S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $F^{m}_{\beta}\uparrow$ $S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $F^{m}_{\beta}\uparrow$
SINet 0.778 0.636 0.043 0.684 0.804 0.718 0.060 0.765
PFNet 0.802 0.666 0.040 0.707 0.826 0.742 0.055 0.781
BGNet 0.833 0.728 0.033 0.759 0.848 0.785 0.046 0.817
FEDER 0.824 0.721 0.032 0.756 0.844 0.785 0.046 0.821
ZoomNet 0.840 0.733 0.030 0.770 0.851 0.782 0.045 0.815
HitNet 0.873 0.810 0.024 0.826 0.873 0.831 0.038 0.851
PRNet 0.875 0.804 0.022 0.826 0.890 0.846 0.032 0.868
CGD 0.874 0.799 0.023 0.819 0.891 0.848 0.032 0.868
CGNet 0.892 0.828 0.019 0.845 0.902 0.863 0.028 0.880
Table 3: Ablation comparison of proposed components. C: semantics consistency module; D: CLIP; E: class-semantic guidance; F: cross-modal multi-head attention; G: multi-level visual collaboration module.
@@
(M) (G) $S_{\alpha}\uparrow$ $E_{\phi}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $M\downarrow$
I Baseline 24.85 7.80 78.91 0.824 0.914 0.792 0.041
II +C 29.08 17.42 44.40 0.872 0.935 0.793 0.023
III +C,D 321.49 78.13 14.18 0.875 0.924 0.791 0.022
IV +C,D,E 327.01 87.34 12.89 0.880 0.928 0.803 0.021
V +C,D,E,F 342.38 88.94 12.68 0.889 0.937 0.818 0.019
VI +C,D,E,G 346.78 90.87 12.48 0.885 0.934 0.809 0.020
VII +C,D,E,F,G 371.33 101.04 11.20 0.890 0.948 0.824 0.018
Qualitative Evaluation. Fig. 5 presents a visual comparison between our proposed CGNet method and other advanced methods. We evaluated challenging camouflaged images featuring different object types under various extreme scenes. These scenes include multiple objects (the first line), structural degradation (second row), occluded scenes (third row), edge similarity (fourth row), and small objects (last row). Such conditions can confound existing COD methods, leading to localization errors, blurred edges, and incomplete segmentation. As observed in Fig. 5, in the first line, other advanced methods struggle to completely segment the two owls due to the high internal similarity of the samples. In second row, the structural semantics of the katydid are disrupted, causing confusion in segmentation for other methods. In third row, heavy occlusion results in interrupted semantic information, preventing other methods from fully segmenting the target. Similarly, In forth and fifth row, complex edge details and small object sizes challenge other methods, resulting in coarse or incomplete segmentation. In contrast to the aforementioned methods, our CGNet approach overcomes these challenges and demonstrates superior predictive performance across these extreme scenes.
5.3 Ablation Study
Impact of Components. To analyze the effect of different components in detail, we evaluate their performance in Tab. 3. In addition, Fig. 6 shows the predicted results for each component. The groups in the table, from I to VII, are incrementally added with one module at a time, where I represents the base network. Our analysis demonstrates that excluding any component leads to a degradation in the performance of the CGNet model. II represents our proposed the class-guided detector. Tab. 3 and Fig. 6 show that CGD achieves excellent performance on COD10K. However, it fails to completely segment targets with a high degree of camouflage. IV performs better than II and III, which indicates the ability of class-Semantic guidance to mitigate domain differences from multiple sources of visual features. However, it still exhibits some false positives and false negatives in complex environments. Compared to I (Baseline), VII (CGNet) improves the performance by 8.3%, which indicates the model progressively focuses on features that are highly relevant to the object class semantics and achieves complete object segmentation under the guidance of class-level textual information. Importantly, incorporating the “D” component significantly increases the model’s parameter count, mainly due to the addition of the CLIP encoder.
@@ 5.3 Ablation Study
Performance on Unseen and Seen Classes. As stated in Sec. 3, we categorize the CamoClass test samples into two sets $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$. This division allows us to evaluate the generalization capability of different methods. Since the distribution of $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$ varies significantly, we rely on results provided by the authors for a fair comparison. As shown in Tab. 2, we evaluate the performance of various COD methods on them. Compared to other COD methods, the results demonstrate that our CGNet achieves superior performance on both sets, particularly on $\mathcal{C}_{unseen}$. This indicates the strong generalization ability of our framework, which benefits from the integration of class-level textual information. To more intuitively show how each method performs on $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$, we provide a comparison of visual results in the supplementary material.

### v2

@@ 4.2 Class-Guided Detector (CGD)
Overview. To improve the performance of COD, we further propose a novel detector CGD. Which leverages class-promt information to accurately segment camouflaged objects and suppress background noise. Given an input $I_{r}$ of size $W\times H$, We first use PvTv2-B2 Wang et al. (2021) as the visual backbone to extract a set of depth features $\{X_{i}\}_{i=0}^{4}$. As shown in Fig. 3, the semantic-rich feature $X_{4}$, the feature $F_{v}$, and the initially fused multi-modal information $F_{m}$ are input into the class-semantic guidance (CSG). The module aims to explore the internal correlations between the backbone features and the other modalities from both shallow and deep layers, thus generating prompt feature $G_{c}$ that embed both class semantic and camouflaged semantic. These features $\{X_{i}\}_{i=1}^{4}$ and $G_{c}$ are then passed to the semantics consistency module to further extract fine-grained features related to the camouflaged object, ultimately leading to an accurate segmentation. The process iteratively accumulates intermediate segmentation results to decouple the final output $\{P_{i}\}_{i=1}^{4}$.
@@
Semantics Consistency Module (SCM). Traditional COD models rely on a visual backbone that extracts camouflage-semantic features containing both foreground and background clues, but this often results in false negatives and false positives. To address these issues, we propose the SCM. By the guidence of class-semantic prompt feature $G_{c}$, the SCM pays more attention to features with high correlation with guided feature, alleviating the interference of irrelevant noise information. Camouflage mainly hides objects by manipulating channel information (e.g., colors), and in rare cases, it also uses spatial positioning to achieve the concealment effect. So we design tow stages to decouple camouflage strategy, the first stage, spatial localization, focuses on capturing spatial information of visual features related to class-semantic prompt feature, while the second stage, channel refinement, explore channel relationship between features and adaptively selects channel information that align with class semantic to suppress noise information interference. Specifically, we utilize two parallel branches to process multi-level features $X_{i}$ and $X_{i+1}$ independently, facilitating their interaction with class semantic knowledge. Each branch incorporates weighted summation and dot product operations, ensuring a comprehensive interaction between the object class-semantic prompt feature and the multi-level features. This interaction ensures a preliminary semantic consistency between features and class prompt feature. Consequently, it produces $G^{\prime}_{{c}}$ and enriched class semantic features $\tilde{X}_{i+1}$ and $\tilde{X}_{i}$. The process is formulated as follows:
$R_{cls}={Softmax}\left({Conv_{3}}\big{(}X_{i}\ +X_{i+1}\big{)}\otimes G_{c}\right)$ (6)
$\hat{X}_{i}=R_{cls}\otimes X_{i},\quad\hat{X}_{i+1}=R_{cls}\otimes X_{i+1}$ (7)
$G^{\prime}_{c}={FC}\left({Concat}\big{(}\hat{X}_{i+1},\hat{X}_{i}\big{)}\right)$ (8)
$\tilde{X}_{i}=G^{\prime}_{c}\ +\hat{X}_{i},\quad\tilde{X}_{i+1}=G^{\prime}_{c}+\hat{X}_{i+1}$ (9)
@@
Here, $i$ denotes the $i$-th operator, ${FC}$ refers to linear projection, while ${Conv_{3}}$ signifies a $3\times 3$ convolution.
@@
FSEL Sun et al. (2024) None PVT 0.885 0.857 0.040 0.942 0.877 0.799 0.021 0.937 0.892 0.852 0.030 0.941 0.916 0.880 0.022 0.958
RISNet Wang et al. (2024) Depth Map PVT 0.870 0.827 0.050 0.922 0.873 0.799 0.025 0.931 0.882 0.834 0.037 0.926 0.908 0.863 0.024 0.951
VSCode Luo et al. (2024) 2D Prompt Swin 0.873 0.820 0.046 0.925 0.869 0.780 0.025 0.931 0.882 0.841 0.032 0.935 - - - -
@@
CGNet (Ours) CamoClass CLIP, PVT 0.896 0.864 0.036 0.947 0.890 0.824 0.018 0.948 0.904 0.869 0.026 0.949 0.931 0.902 0.017 0.972
To enhance the feature expression and focus on class-relevant information, we propose a channel-adaptive refinement mechanism. Initially, features $\tilde{X}_{i+1}$ and $\tilde{X}_{i}$ are concatenated along the channel dimension to form a unified feature representation $X_{{cot}}$. This representation is then partitioned into $j$ subgroups $\{g_{1},g_{2},\dots,g_{j}\}$ along the channel dimension, where $j=4$. For each subgroup $g_{j}$, its concern weight $w_{j}$ is calculated through the linear layer and sigmoid activation function. These weights modulate the importance of each subgroup by element-wise multiplication, emphasizing subgroups that are consistent with the semantics of the object class. The weighted subgroups are subsequently fused along the channel dimension to produce the refined feature representation $F^{i}_{{s}}$, which highlights the critical information associated with the object class. The process is defined as follows:
$X_{{cot}}=Conv_{3}({Concat}(\tilde{X}_{i},\tilde{X}_{i+1}))$ (10)
$\{g_{1},g_{2},\dots,g_{j}\}={Split}(X_{{cot}})$ (11)
$w_{j}={Gate}(g_{j})$ (12)
$g^{\prime}_{j}=Conv_{3}(x_{{cot}}\otimes w_{j})$ (13)
$F^{i}_{{s}}=Conv_{3}({Concat}(g^{\prime}_{1},g^{\prime}_{2},\dots,g^{\prime}_{j}))$ (14)
where $Gate$ refers to tow linear projection and a sigmoid.
@@ 4.3 Loss Function
Since our ultimate goal is to generate a binary foreground graph, a structured loss function $\mathcal{L}_{seg}=\mathcal{L}_{{bce}}+\mathcal{L}_{{iou}}$ consisting of BCE loss ($\mathcal{L}_{{bce}}$) and IoU loss ($\mathcal{L}_{{iou}}$) can be used in the optimization process Wei et al. (2020). Specifically, the loss function can be expressed as:
$\mathcal{L}=\sum_{i=1}^{4}\mathcal{L}_{{seg}}(P_{i},{G})+\mathcal{L}_{{seg}}(F_{v},{G})+\mathcal{L}_{{seg}}(F_{m},{G})$ (15)
where ${F_{m},F_{v},P_{4},P_{3},P_{2},P_{1}}$ represent the predictions generated by our model, and ${G}$ refers to the ground truth.
@@ 5.1 Implementation Details
Data Settings. We used proposed CamoClass dataset and traditional employ the traditional train paradigm to evaluate our approach. The training set consists of CAMO Le et al. (2019) and COD10K Fan et al. (2022), with a total of 4040 samples and their class-level textual information. The test set consists of test samples from four benchmark datasets, totaling 8615 samples and their class-level textual information.
Evaluation Metrics. In this work, we use five widely recognized and standard metrics to assess the performance of our method, including S-measure ($S_{\alpha}$an et al. (2017), E-measure ($E^{m}_{\phi}$an et al. (2018), weighted F-measure ($F_{\beta}^{\omega}$argolin et al. (2014), average F-measure ($F^{m}_{\beta}$) and mean absolute error ($M$).
Training Settings. To reduce computational load and leverage CLIP’s pre-trained knowledge Xu et al. (2023), we keep the CLIP model frozen in our experiments, using the ViT-L@336. while the parameters of the remaining network components were learnable. Following the setup in previous studies Pang et al. (2022); Fan et al. (2022), We use the Adam optimizer with as learning rate of $1\times 10^{-4}$ to optimize the network parameters, with a batch size of 10 and a total of 200 epochs. We apply various data augmentation techniques to the training data, including random cropping, horizontal flipping, and color jittering. Specially, during both the training and inference phases, the image input size for the CLIP model was $336\times 336$ pixels, while the input size for the class-guided detector was $448\times 448$ pixels. All experiments are conducted on a server equipped with an NVIDIA RTX 3090 (with 24GB memory).
@@ 5.2 Comparison with State-of-the-Art Methods
Quantitative Evaluation. As shown in Tab. 1, CGNet consistently surpasses competing methods across all datasets, showing a significant and substantial performance advantage. And CGNet are on average 5.5% better than the second best method. Due to CRIS, CGFormer, and ASDA are not specifically designed for the COD task, their performance on the four datasets is poor. Moreover, we also present the results of detectors optimized under the class prompt generator (CPG) and proposed CamoClass. In Tab. 1, CPG improve other detectors by 19.73% (SINet), 13.3% (BGNet), 12.68% (PFNet), 4.43% (FPSNet), and increase our class-guided detector by 6.24% (PVT), verifying that CPG serve as effective plug-and-play module for performance improvement.
Table 2: Comparison of methods on $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$.
Methods $\mathcal{C}_{seen}$ (6109 images) $\mathcal{C}_{uneen}$ (364 images)
$S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $E_{m}\uparrow$ $S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $E_{m}\uparrow$
SINet 0.797 0.693 0.054 0.868 0.786 0.680 0.051 0.858
PFNet 0.820 0.719 0.050 0.884 0.801 0.694 0.051 0.868
BGNet 0.845 0.768 0.042 0.905 0.827 0.741 0.043 0.873
ZoomNet 0.849 0.768 0.039 0.894 0.826 0.733 0.043 0.875
FEDER 0.840 0.767 0.041 0.905 0.813 0.725 0.045 0.873
HitNet 0.875 0.827 0.033 0.929 0.856 0.797 0.035 0.916
RISNet 0.880 0.825 0.033 0.928 0.860 0.794 0.038 0.910
PRNet 0.887 0.834 0.029 0.942 0.864 0.801 0.035 0.912
FSEL 0.889 0.838 0.027 0.943 0.867 0.807 0.032 0.921
CGD 0.887 0.831 0.029 0.937 0.861 0.796 0.035 0.914
CGNet 0.900 0.854 0.024 0.954 0.882 0.819 0.029 0.926
Qualitative Evaluation. Fig. 5 presents a visual comparison between our proposed CGNet method and other advanced methods. We evaluated challenging camouflaged images featuring different object types under various extreme scenes. These scenes include multiple objects (the first line), structural degradation (second row), occluded scenes (third row), edge similarity (fourth row), and small objects (last row). Such conditions can confound existing COD methods, leading to localization errors, blurred edges, and incomplete segmentation. As observed in Fig. 5, in the first line, other advanced methods struggle to completely segment the two owls due to the high internal similarity of the samples. In second row, the structural semantics of the katydid are disrupted, causing confusion in segmentation for other methods. In third row, heavy occlusion results in interrupted semantic information, preventing other methods from fully segmenting the object. Similarly, In forth and fifth row, complex edge details and small object sizes challenge other methods, resulting in coarse or incomplete segmentation. In contrast to the aforementioned methods, our CGNet approach overcomes these challenges and demonstrates superior predictive performance across these extreme scenes.
5.3 Ablation Study
Table 3: Ablation comparison of proposed components. Baseline: PVT; C: semantics consistency module; D: CLIP; E: class-semantic guidance; F: cross-modal multi-head attention; G: multi-level visual collaboration module.
@@ 5.3 Ablation Study
(M) (G) $S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $E_{\phi}^{m}\uparrow$
I Baseline 24.85 7.80 78.91 0.824 0.792 0.041 0.914
II +C 29.08 17.42 44.40 0.872 0.793 0.023 0.935
III +C,D 321.49 78.13 14.18 0.875 0.791 0.022 0.924
IV +C,D,E 327.01 87.34 12.89 0.880 0.803 0.021 0.928
V +C,D,E,F 342.38 88.94 12.68 0.889 0.818 0.019 0.937
VI +C,D,E,G 346.78 90.87 12.48 0.885 0.809 0.020 0.934
VII +C,D,E,F,G 371.33 101.04 11.20 0.890 0.824 0.018 0.948
Impact of Components. To analyze the effect of different components in detail, we evaluate their performance in Tab. 3. In addition, Fig. 6 shows the predicted results for each component. The groups in the table, from I to VII, are incrementally added with one module at a time, where I represents the base network. Our analysis demonstrates that excluding any component leads to a degradation in the performance of the CGNet model. II represents our proposed the class-guided detector. Tab. 3 and Fig. 6 show that CGD achieves excellent performance on COD10K. However, it fails to completely segment objects with a high degree of camouflage. IV performs better than II and III, which indicates the ability of class-semantic guidance to mitigate domain differences from multiple sources of visual features. However, it still exhibits some false positives and false negatives in complex environments. Compared to I (Baseline), VII (CGNet) improves the performance by 8.3%, which indicates the model progressively focuses on features that are highly relevant to the object class semantics and achieves complete object segmentation under the guidance of class-level textual information. Particularly, incorporating the “D” component significantly increases the model’s parameter count, mainly due to the addition of the CLIP encoder.
@@
Performance on $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$. As stated in Sec. 3, we categorize the CamoClass test samples into two sets $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$. This division allows us to evaluate the generalization capability of different methods. Since the distribution of $\mathcal{C}_{seen}$ and $\mathcal{C}_{unseen}$ varies significantly, we rely on results provided by the authors for a fair comparison. As shown in Tab. 2, we evaluate the performance of various COD methods on them. Compared to other COD methods, the results demonstrate that our CGNet achieves superior performance on both sets, particularly on $\mathcal{C}_{unseen}$. This indicates that our framework demonstrates a certain degree of generalization ability and zero-shot capability, benefiting from the integration of class-level textual information. Importantly, we provide more detailed analysis and comparison of visual results in the supplementary material.
Table 4: Comparison of methods on hard and normal subsets.
Methods $Hard$ (6223 images) $Normal$ (250 images)
$S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $E_{m}\uparrow$ $S_{m}\uparrow$ $F^{\omega}_{\beta}\uparrow$ $\mathcal{M}\downarrow$ $E_{m}\uparrow$
SINet 0.793 0.686 0.054 0.876 0.871 0.841 0.045 0.884
PFNet 0.816 0.712 0.050 0.892 0.887 0.856 0.040 0.897
BGNet 0.841 0.762 0.042 0.912 0.900 0.881 0.035 0.916
ZoomNet 0.845 0.762 0.040 0.910 0.907 0.884 0.032 0.918
FEDER 0.836 0.760 0.041 0.908 0.896 0.880 0.035 0.914
HitNet 0.872 0.822 0.033 0.930 0.912 0.902 0.028 0.926
RISNet 0.877 0.819 0.034 0.931 0.921 0.909 0.027 0.933
PRNet 0.884 0.829 0.029 0.942 0.924 0.909 0.025 0.933
FSEL 0.886 0.833 0.028 0.946 0.923 0.912 0.025 0.936
CGD (ours) 0.884 0.826 0.030 0.941 0.921 0.904 0.028 0.929
CGNet (ours) 0.898 0.849 0.025 0.952 0.936 0.925 0.022 0.946
Performance in Different Camouflaged Scenes. In Tab. 4, we evaluate the performance of our CGNet and the existing methods in different camouflage degree scenes by splitting the CamoClass test samples based on the detection results of CGNet. Samples with $S_{m}\geq 0.9$ are categorized as $Normal$, while those with $S_{m}<0.9$ are labeled as $Hard$. The relevant analysis is detailed in the supplementary materials.